# https://github.com/krulllab/COSDD

### v2

@@ 5.5 Ablation study - noise reconstruction
Figure 5: A noisy image from the Convallaria A dataset was encoded and decoded to produce a reconstructed observation and an artificial noise sample. We examine this sampled noise to check that its spatial autocorrelation and signal dependence match those of the real noise, indicating that the AR decoder accurately models the noise.
If the VAE’s decoder is modeling only the noise component of images, encoding an image with the VAE and sampling from the AR decoder should yield an image with the same underlying signal but a different random sample of noise. If the decoder’s model of the noise is accurate, the sampled noise should exhibit the same autocorrelation and signal-dependence characteristics as the noise in the original image. An investigation into this is reported in Fig. 5, where a noisy image from the Convallaria A dataset was encoded by a trained VAE, then a latent variable was sampled and used by the VAE’s AR decoder to sample a reconstruction of the original image. The reconstructed noise exhibits spatial autocorrelation and signal dependence very similar to the real noise, indicating that our AR decoder has learnt an accurate model of the noise.
@@ 6 Conclusion
We proposed an unsupervised VAE-based denoising algorithm for noise that is correlated along rows or columns of pixels and is signal-dependent. By engineering the receptive field of the AR decoder, the VAE’s latent variables are encouraged to represent the signal content of an image while discarding the noise. We then presented a novel signal decoder that is trained to map this latent variable into an estimate of the clean image. The algorithm outperforms both the self-supervised denoiser SN2V [2] and the unsupervised denoiser HDN36 [35].
Our method is suited to noise with correlations that run parallel to the axes of the image. Such noise commonly occurs in microscopy. Often, users in labs are unable to find suitable unsupervised methods to remove such noise. We release our code as open source and strongly believe that the scientific imaging community will apply and adapt our methods in a variety of applications.
While we have achieved our results using simple 1-dimensional receptive fields, some imaging modalities produce noise that is correlated in multiple directions, therefore requiring a differently shaped receptive field. We found that extending the receptive field to cover both a row and column of pixels allows the decoder to model some aspects of the signal, making the technique unsuitable for removing noise correlated in two dimensions. However, we believe that techniques other than shaping the AR decoder receptive field could be discovered to limit the decoder’s modelling capabilities. We hope that future work will further improve the theoretical understanding of the method and allow us to utilize its full potential.
Acknowledgements We would like to thank Ales Leonardis (University of Birmingham) for a helpful discussion on the writing of the paper. The computations described in this paper were performed using the University of Birmingham’s BlueBEAR HPC service, which provides a High Performance Computing service to the University’s research community. See http://www.birmingham.ac.uk/bear for more details.

### v3

@@ 5.3 Ablation Study - Noise Reconstruction
Figure 5: A noisy image from the Convallaria A dataset was encoded and decoded to produce a reconstructed observation and an artificial noise sample. The sampled noise has spatial autocorrelation and signal dependence that match those of the real noise, indicating that the AR decoder has learnt an accurate noise model.
If the VAE’s AR decoder is modeling only the noise component of images, encoding an image with the VAE and sampling from the AR decoder should yield an image with the same underlying signal but a different random sample of noise. If the decoder’s model of the noise is accurate, the sampled noise should exhibit the same autocorrelation and signal-dependence characteristics as the noise in the original image. An investigation into this is reported in Fig. 5, where a noisy image from the Convallaria A dataset was encoded by a trained VAE, then a latent variable was sampled and used by the VAE’s AR decoder to sample a reconstruction of the original image. The reconstructed noise exhibits spatial autocorrelation and signal dependence very similar to the real noise, indicating that our AR decoder has learnt an accurate model of the noise.
@@ 6 Conclusion
We proposed an unsupervised VAE-based denoising algorithm for signal-dependent noise that is correlated along rows or columns of pixels. It is trained without any clean data or a noise model. By engineering the receptive field of the AR decoder, the VAE’s latent variables are encouraged to represent the signal content of an image while discarding the noise. We then presented a novel signal decoder that is trained to map this latent variable to an estimate of the clean image. The algorithm outperforms existing self- and unsupervised denoisers.
Our method is suited to noise with correlations that run parallel to the image axes. This commonly occurs in microscopy and lab users often cannot find suitable methods to remove it. We release our code open source and strongly believe that the scientific imaging community will apply and adapt our methods in a variety of applications.
While we have achieved our results using 1-dimensional receptive fields, some imaging modalities produce noise that is correlated in multiple directions, therefore requiring a differently shaped receptive field. We found that extending the receptive field to cover both a row and a column of pixels allows the AR decoder to model some aspects of the signal, making the technique unsuitable for removing noise correlated in two dimensions. However, we believe that techniques other than shaping the AR decoder’s receptive field could be discovered to limit it’s modeling capabilities. Furthermore, the method is limited to zero-centered noise by the signal decoder. This precludes, e.g., removing salt-and-pepper noise or restoring images with saturated pixel intensities. A direction for future work could therefore be alternative methods for mapping latent variables back into image space. We hope that future work will further improve the theoretical understanding of the method and allow us to utilize its full potential.
Acknowledgements We thank Aleš Leonardis (University of Birmingham) for a helpful discussion on the writing of this paper. Computations in this paper were performed using the University of Birmingham’s BlueBEAR HPC service, which provides a High Performance Computing service to the University’s research community. See http://www.birmingham.ac.uk/bear for more details.